OpenStudy (anonymous):

How many grams of magnesium oxide can be produced when 125 g Mg react with 102 g O2?

OpenStudy (anonymous):

i already balenced the equation and got: 2MgO + O2 yield 2MgO

OpenStudy (jfraser):

Find the moles of each reactant that are available\[mol Mg = 125g Mg * (\frac{1mol Mg}{24.3g Mg})\]\[mol O{_2} = 102g O{_2} * (\frac{1mol O{_2}}{32g O{_2}})\] The reaction requires a 2:1 ratio. One of those 2 is in short supply.

OpenStudy (anonymous):

so is the oxygen short?

OpenStudy (anonymous):

??

OpenStudy (jfraser):

There are 5.14 moles of Mg available, and 3.19 moles of O2 available. The ratio required is 2Mg :1O2

OpenStudy (anonymous):

so how do i find out the answer to the qustion?

OpenStudy (jfraser):

Does 5.14/3.19 = 2/1?

OpenStudy (anonymous):

it equals 1.611

OpenStudy (jfraser):

If you were going to use all 3.19 moles of O2, there would have to be twice as much Mg (6.28moles). There isn't that much available, there's only 5.14moles. The Mg will run out first. Because of that, the Mg is the limiting reactant, it will limit the amount of product you can form.

OpenStudy (anonymous):

i understand tht but how do i find out how much magnesium i can make

OpenStudy (jfraser):

Use the coefficients of the balanced reaction: 2 Mg will make 2MgO